OpenStudy (anonymous):

A single die is rolled, then a card is drawn from a deck containing only face cards (i.e. only Jack, Queen, King of each suit). Finally, a coin is flipped. You win the game if you toss tails, draw a heart, and roll an even number. Find the probability that you do not win.

OpenStudy (anonymous):

I got \[=\frac{ 1*4*3*3 }{ 144 }\] \[=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

P(lose) = 1 - P(win) =1 - (1/2) (3/12) (3/6) = 15/16